# A Learning-Based Framework for Collision-Free Motion Planning

### v1

@@ Abstract
Index Terms:
Motion planning, robot learning, deep learning, circular fields, potential fields, predictive agents, Bayesian optimization, manipulability, obstacle avoidance, robot perception
@@ VI-A Simulation
Simulation results for two different scenarios are shown in Figures 3 and 4. The first and last frames of each trajectory are illustrated.
@@ VI-B Robot experiments
Figures 6 and 7 show the results of horizontal and vertical obstacle avoidance tasks, respectively
Figure 7: First, intermediate, and last frames of the simulation. This sequence illustrates the robot’s motion progression from initialization to task completion.
Figure 8: First, intermediate, and last frames of the simulation. This sequence illustrates the robot’s motion progression from initialization to task completion.
@@ VII Discussion and Conclusion
In this paper, we propose a learning-based approach that infers the optimal parameter vector for a circular-field-based motion planner. Furthermore, we demonstrated the insufficiency of the original planner when predefined gains are used.

### v2

@@ VI-A Simulation
Simulation results for two different scenarios are shown in Figures 5 and 6. The first and last frames of each trajectory are illustrated.
@@ VI-B Robot experiments
Figures 7 and 8 show the results of horizontal and vertical obstacle avoidance tasks, respectively
Figure 7: First, intermediate, and last frames of the horizontal obstacle avoidance experiment. This sequence illustrates the robot’s motion progression from initialization to task completion.
Figure 8: First, intermediate, and last frames of the vertical obstacle avoidance experiment. This sequence illustrates the robot’s motion progression from initialization to task completion.
@@ VII Discussion and Conclusion
In this paper, we propose a learning-based approach that infers the optimal parameter vector for a circular-field-based motion planner. Furthermore, we demonstrate the insufficiency of the original planner when predefined gains are used.